OpenStudy (jennychan12):

how to integrate sint^2)

OpenStudy (anonymous):

is it sin(t^2) or sin^2(t)

OpenStudy (jennychan12):

\[\int\limits_{0}^{x^3} \sin(t^2)dt\]

OpenStudy (anonymous):

okay so do you know the anti-derivative of sine?

OpenStudy (jennychan12):

-cos

OpenStudy (anonymous):

correct

OpenStudy (jennychan12):

i'm just confused with the t^2 in the inside

OpenStudy (anonymous):

well if you where doing the derivative of -cos(t^2) what would you get?

OpenStudy (jennychan12):

2tsin(t^2) ?

OpenStudy (anonymous):

right, but as we see in the integral there isnt a 2t, so therefore we must divide by 2t so cancel this out. Does that make sense?

OpenStudy (anonymous):

Have you done integration using substitution?

OpenStudy (jennychan12):

yes.

OpenStudy (anonymous):

u should make a substitution. u = t^2

OpenStudy (anonymous):

then change the dt to du and change the limits and u then should just have to integrate sin u

OpenStudy (jennychan12):

i origninally thought the antiderivative was \[-\frac{ t^3 }{ 3 } \cos(t^2)\]

OpenStudy (anonymous):

shall I show u my method?

OpenStudy (jennychan12):

sure.

OpenStudy (anonymous):